# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

### v1

@@ Appendix B Choice of residuals
This appendix briefly discusses the choice of residuals $\{a_{t}\}$ that are used in the proof of Proposition 2.2(a) through Lemma 2.1.
In the setup of Proposition 2.2(a), it is straightforward to show that if $\{a_{t}\}$ is a nonnegative sequence of scalars such that
@@
then repeatedly applying Lemma 2.1 with $a=a_{t}$ yields
@@ Appendix C Parameter choices
$=\frac{1-L^{-2}}{1-(\xi-1)/\xi}=\xi\left(1-\frac{1}{L^{2}}\right).\qed$
@@
for every $\rho>0$. If $\lambda=\overline{\lambda}(m)$, $\sigma\geq\overline{\sigma}_{\varepsilon}(m)$, and $\ell\geq\overline{\ell}(m)$, then
@@
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $b$, $C$, $E$, $\ell$, $\lambda$, and $T$ be as in Theorem 4.5. If
$\lambda\leq\min\left\{\frac{1}{\sqrt{E}},\frac{1}{2(m+M)}\right\},$
then we have
@@
$\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[\frac{1}{E}+\theta_{L_{\lambda}}(\ell)\right]\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[1+\theta_{L_{\lambda}}(\ell)\right].\qed$

### v3

@@ Appendix B Choice of residuals
This appendix briefly discusses the choice of residuals $\{a_{t}\}$ that are used in the proof of Proposition A.5(a) and Lemma A.2.
In the setup of Proposition A.5(a), it is straightforward to show that if $\{a_{t}\}$ is a nonnegative sequence of scalars such that
@@
then repeatedly applying Lemma A.2 with $a=a_{t}$ yields
@@ Appendix C Parameter choices
$=\frac{1-L^{-2}}{1-(\xi-1)/\xi}=\xi\left(1-\frac{1}{L^{2}}\right).$
@@
for every $\rho>0$. If $\lambda=\overline{\lambda}(m)$, $\sigma\geq\overline{\sigma}_{\varepsilon}(m)$, $\ell\geq\overline{\ell}(m)$, and no gradient clipping is performed, then
@@
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $b$, $C$, $E$, $\ell$, $\lambda$, $\sigma$, and $T$ be as in Theorem 2.3. If
$\lambda\leq\min\left\{\frac{1}{\sqrt{E}},\frac{1}{2(m+M)}\right\}$
and no gradient clipping is performed, then we have
@@
$\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[\frac{1}{E}+\theta_{L_{\lambda}}(\ell)\right]\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[1+\theta_{L_{\lambda}}(\ell)\right].$
Appendix D Limitations of Poisson sampling in practice
This appendix discussing the computational limitation of implementing Poisson sampling in practice. It is primarily concerned with the large-scale setting where datasets may be on the order of hundreds of millions of examples.
Data access. Implementations of Poisson sampling, e.g., Opacus [29], typically employ a pseudorandom number generator to (i) randomly sample a collection of indices from zero to N1N-1, where $N$ is the size of the dataset and (ii) map these indices to corresponding examples in the dataset to generate a batch of examples. In order for (ii) to be efficient, many libraries need fast random access to the dataset which is difficult to do without loading the entire dataset into RAM (as reading data from disk can be orders of magnitude more expensive). In contrast, cyclic traversal of batches only requires (relatively small) fixed blocks of the dataset to be loaded into memory for every batch and need not perform a matching of indices (such as in (i) above) to data.
Variable batch size. Independent of the access speed of the dataset examples, Poisson sampling also generates batches of random sizes, which are typically inconvenient to handle in deep learning systems [14]. For example, popular just-in-time compilation-based machine learning libraries such as JAX, PyTorch/Opacus, and TensorFlow may need to retrace their computation graph at every training step as the batch size cannot be statically inferred or kept constant. Additionally, optimizing training workloads on hardware accelerators such as graphical processing units (GPUs) and tensor processing units (TPUs) becomes difficult as (i) they require any in-device data to have fixed sizes and (ii) any input data generated by Poisson sampling will have variable sizes due to the effect of variable batch sizes. In contrast, the cyclic traversal of batches will always generate fixed batch sizes and, consequently, will not suffer from the above issues.